OpenStudy (anonymous):

Choose the statement that correctly lists equivalent ratios. Select one of the options below as your answer: A. 1 : 6 = 20 : 120 = 5 : 30 B. 1 : 6 = 4 : 18 = 6 : 36 C. 1 : 8 = 20 : 120 = 3 : 24 D. 1 : 8 = 3 : 32 = 5 : 48

OpenStudy (anonymous):

A

OpenStudy (anonymous):

how'd you do it?

OpenStudy (anonymous):

use the 1 as a basepoint, 20/1=20 so 6*20 we get 120 similarly 5/1=5 so 6*5=30

OpenStudy (anonymous):

for the others it does not work out

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem